OpenStudy (anonymous):

what is the equation in this word problem? Suzanne is going to rent a car while she is out of town. One car rental company offers a flat rate of $35 per day plus $0.10 per mile. Another car rental company offers the same car for $25 per day plus $0.25 per mile. She will need the car for 5 days. How many miles would she need to drive for the first rental company to be a better deal?

OpenStudy (ybarrap):

r1 = 35*d + .10*m, rental car company 1 r2 = 25*d + .25m, rental car company 2 r1 will be a better deal when when r2 > r1 r2 > r1 when 35*5 + .10*m > 25*5 + .25m 5(35-25) > (.25 - .10)m 5*10 > (.15)m 50/.15 > m m > 50/.15 = 33 1/3 Ans. When 33 1/3 miles have been driven, rental car company r1 will be a better deal.